The wavelength associated with a golf ball weighing 200g and moving at a speed of 5 m/h is of the order
a. 10-10 m
b. 10-20 m
c. 10-30 m
d. 10-40 m

Grade:upto college level

3 Answers

Navjyot Kalra
9 years ago
Using the de-Broglie’s relationship : λ = h/mv = 〖6.625 × 10〗^(-34)/(0.2 × 5/(60 × 60)) = 2.3 × 10-30 m
Bhupendra Verma
5 years ago
Wavelength=h/mv,
                     =6.6*10-34 j-s/ 200*10-3kg*5/3600
                     =3.3*10-31*36/5
                     =118.8*10-31/5
                     =23*10-31
                    =2.3*10-31 (ans)
                    =10-30 (ans)
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Using the de-Broglie’s relationship :
λ = h/mv
= 〖6.625 × 10〗^(-34)/(0.2 × 5/(60 × 60))
= 2.3 × 10^-30 m

Thanks and Regards